OpenStudy (anonymous):

Solve using either Elimination or Substitution. Want to check against my answer. y=1/2x -6 2x+6y =19

OpenStudy (anonymous):

x=11 y=-1/2 substitution elimination

OpenStudy (anonymous):

It cannot be solve by elimination because it is dependent.

OpenStudy (anonymous):

I like using elimination because it is much faster. just write everything in the form of x+y=c where c is a constant such as 2 and then use elimination.

OpenStudy (anonymous):

I'll set this up for you \[2x+6y =19\] \[1/2 x-y=6\] Now, what can you do to \[1/2 x\] to make it look like \[2x\]? Because the point of elimination, is that either the x cancel each other or the y cancel each other.

OpenStudy (anonymous):

@beeqay, Thanks for your help, I appreciate it but recheck, isn't Substitution the better way of solving it?

OpenStudy (anonymous):

It totally depends on personal opinion, if you feel that it is easier for you then it is. I personally think elimination is easier, if you multiply the equation with \[1/2 x\] by 6 then you will have 3x-6y=36. So now you can solve this by elimination easily since the 6y cancels out with the 6y in the other equation. and you have 5x=55.